What is the probability that a child, picked at a random, has a greater than 116 cm?


1
Expert's answer
2021-04-15T07:52:18-0400


Average height "\mu =110cm"

Standard deviation "\sigma=6 cm"


Z score is = "\dfrac{x-\mu}{\sigma}"


      "=\dfrac{116-110}{6}=1"


"P(X>116)=P(z>1)=0.1587"